OpenStudy (anonymous):

Mia deposits either $5 or $3.50 into her savings account. She made 14 deposits equaling a total of $59.50. How many times did she deposit $5 and how many times did she deposit $3.50? Mia made deposits of $5 and deposits of $3.50.

OpenStudy (anonymous):

@AriPotta

OpenStudy (dtan5457):

are you still there?

OpenStudy (anonymous):

yes

OpenStudy (dtan5457):

give me a few min

OpenStudy (anonymous):

ok

OpenStudy (dtan5457):

ok i got it, let me explain

OpenStudy (anonymous):

ok

OpenStudy (dtan5457):

let x represent the amount of deposits worth 5 dollars let y represent the amount of deposits worth $3.50 make a system x+y=14<<all the deposits there are 5x+3.50y=59.50 multiply x+y=14 by -5, so you can solve for a variable you get 5x+3.50y=59.50 -5x-5y=-70 you add that up, and you get -1.50y=-10.50 find y by dividing it by -1.50 you get y=7 if y=7 and 14-7=7 meaning x also equals 7. this is right as 5 times 7= 35 3.50 times 7=24.5 35+24.5=59.50 >>>> seven $5 dollar deposits >>>> seven $3.50 deposits

OpenStudy (anonymous):

thank you. word problems are very tricky